Ten mens can finish a work in 15days.15womens can finish a same work in 12days. How many days are required to complete the same work If,both men &women are combine together?






Answer by ptaylor(2198)

Let x=time it takes if both are working together
10 men work at the rate of 1/15 of the job per day
15 women works at the rate of 1/12 of the job per day
Working together, 10 men and 15 women work at the rate of (1/15)+(1/12)=(4/60)+(5/60)=9/60=3/20 of the job per day
Now our equation to solve is:
(3/20)*x=1 (1 job, that is)
3x=20
x=6 2/3 days
CK
In 6 2/3 =20/3 days, 10 mens can do (20/3)(1/15) =4/9 of the job
In 20/3 days 15 women can do (20/3)(1/12)=5/9 of the job
5/9 + 4/9= THE JOB!!!
Hope this helps----ptaylor